# Reinforcement Inference: Leveraging Uncertainty for Self-Correcting Language Model Reasoning

### v1

@@ I Introduction
Despite the remarkable performance, LLMs can produce confidently incorrect answers. A well-known challenge is that models often lack calibrated confidence – they may not recognize when their answers are likely wrong [1, 2]. This can lead to misleading outputs in high-stakes settings (for example, an LLM hallucinating a plausible-sounding but false explanation with full certainty [3, 4]). If an AI could know when it is unsure, it might mitigate errors by double-checking or deferring answers. Indeed, prior studies on tasks like multiple-choice QA have noted that an LLM’s softmax probabilities carry information about correctness [5, 2]. Intuitively, a model that assigns one option nearly 100% probability is likely confident (and often correct), whereas if it spreads probability roughly uniformly across all options, it is essentially guessing [5, 2]. Recent evaluations confirm that wrong answers tend to have lower maximum softmax probability (MSP) than correct answers in well-trained models [2, 6]. In short, uncertainty correlates with mistakes.
How can we use this insight to improve model reasoning? One approach is to have the model abstain or refuse to answer when uncertainty is high [7]. Refusal might be safer in open-ended applications, but in many scenarios we do want the model to attempt an answer – just with more care if it was initially unsure. Humans do this naturally: if unsure of an answer, a person might pause and double-check their reasoning or try a different approach before responding. This inspires the question: Can an LLM similarly “think twice” when it hesitates, to produce a more accurate answer, all at inference time?
@@
In this paper, we propose an inference-time framework called Reinforcement Inference to realize this idea. The core concept is simple: the model’s own inferred uncertainty and MSP are used as a trigger for a second reasoning attempt. If the model’s first answer had high entropy – indicating confusion or low confidence – we prompt the model again with a special instruction that it was uncertain and should reason more thoroughly. Crucially, we do not reveal the correct answer or provide any external feedback; the model must rely on its existing knowledge, but is now explicitly guided to be careful. This procedure requires no additional training or parameter updates. It functions as a form of test-time control or meta-cognition for the model, akin to a student checking their work upon sensing uncertainty. As we will show, even this light-touch intervention can yield significant accuracy improvements on challenging questions that the model initially struggled with.
Before diving into technical details, we emphasize the broader motivation of this work: modern LLM capability is often not fully realized by the default inference protocol. In many deployed settings—especially in professional and industrial workflows that require deterministic behavior—models are typically used with greedy decoding and a single forward pass. This “one-shot, greedy” regime can systematically under-estimate what a fixed model is actually capable of: many failures are not due to missing knowledge, but due to shallow first attempts, premature commitment, or unresolved internal ambiguity. Reinforcement Inference provides a practical way to probe and unlock a model’s latent capability horizon without any retraining. The large accuracy jump we observe is therefore not merely a reprompting trick; it is evidence that a substantial portion of the model’s reasoning capacity is available at inference time but remains untapped under the standard one-pass protocol.
Methodologically, our key idea is to treat uncertainty as a first-class control signal during inference. Entropy and MSP expose when the model is effectively “guessing” under its own posterior, and we convert this signal into an adaptive compute policy: allocate a second reasoning pass only when the model itself indicates hesitation. This yields an efficiency–accuracy paradigm suited to real systems: predictable latency on easy cases, and deliberate reasoning only when needed.
More broadly, we argue that entropy-aware inference suggests an entropy-aware training direction. For professional models where correctness and determinism matter, training objectives and decoding policies should explicitly constrain the relationship between correctness and confidence (e.g., penalizing high-confidence errors and encouraging calibrated decisiveness). Our work offers a concrete blueprint for such future designs: (i) use entropy/MSP to define when additional reasoning is warranted, (ii) quantify the resulting “capability gap” between one-pass greedy inference and uncertainty-conditioned deliberation, and (iii) use this gap as a diagnostic signal for both decoding and model optimization. In this sense, Reinforcement Inference is a step toward a broader paradigm in which uncertainty is not only measured, but also systematically acted upon.
@@ III-A Confidence and Uncertainty in LLMs
There is growing interest in measuring how well an LLM’s predicted probabilities reflect its actual likelihood of being correct [2, 15], including recent work on whether LLMs can reason explicitly about their own confidence [44]. Maximum softmax probability (MSP) has been used as a simple proxy for model confidence, with studies showing that lower MSP correlates with wrong answers and even with out-of-distribution inputs in both vision and NLP models [2, 15]. For instance, Plaut et al. (2024) provide strong evidence that in multiple-choice QA, an LLM’s softmax distribution can predict whether its answer is correct [16, 17]. They leveraged this by allowing the model to abstain on low-confidence answers, improving overall accuracy by avoiding some mistakes [46, 18, 17]. Other works explore more complex uncertainty quantification methods (Bayesian approaches, ensembles, or probing internal states [17, 19]) and calibrating models via fine-tuning to output better confidence estimates [20, 21]. Our work aligns with these efforts in that we use the model’s own probability distribution as an uncertainty signal. However, instead of training the model to be calibrated or to abstain, we use uncertainty to trigger a self-correction process at inference time.
@@ III-B Chain-of-Thought and Self-Consistency
Prompting strategies like chain-of-thought (CoT) encourage models to generate intermediate reasoning steps, which often boosts performance on complex tasks [8]. Building on CoT, the self-consistency approach samples multiple independent reasoning paths and then selects the most common answer among them [8, 9]. Wang et al. (2023) showed that self-consistency significantly improves accuracy on math and commonsense questions by effectively ensembling reasoning outcomes [9]. While powerful, self-consistency incurs substantial computational cost (many model forward passes per query) and always uses multiple attempts even if the model was confident on the first try. In contrast, our method performs a second pass selectively – only for cases where the first pass was uncertain. Thus, Reinforcement Inference can be seen as a targeted, single-step variant of reasoning augmentation. It resembles asking the model to “try a different line of reasoning” but only when warranted by low confidence, saving computation on easier queries.
@@ III-C Verifier Models and Answer Re-Ranking
Another line of work improves answer correctness by introducing a verification stage. One paradigm is to train a separate verifier classifier that judges whether the model’s answer is correct (sometimes using explanations or retrieved evidence), and then use that to accept or reject answers [10, 11]. Other approaches prompt the model itself to check its answer. For example, Chain-of-Verification (CoVe) has the model generate an initial answer, then pose follow-up verification questions to itself, answer those, and finally produce a revised answer [22, 23]. This multi-step deliberation can reduce factual hallucinations [24, 25]. Similarly, self-critique techniques have the model reflect on and potentially correct its first answer by analyzing its own output for errors [26, 27]. These verification-based methods improve reliability, but they often rely on complex prompt programs or fine-tuned models for critique. Our approach is much simpler: we do not generate explicit verification questions or use a separate model; we just give a single additional prompt reminding the model to be careful. In effect, Reinforcement Inference is a lightweight self-reflection step triggered by uncertainty, rather than a full-fledged self-dialogue or learned verifier.
@@ III-D Supervised Retraining for Reasoning
A straightforward way to improve a model’s accuracy or calibration is additional training. Fine-tuning on high-quality reasoning data (e.g. with CoT exemplars or human explanations) can make a model more reliable on tricky problems. Reinforcement Learning from Human Feedback (RLHF) is also widely used to align models to preferred answers and avoid obvious mistakes. For example, recent work has fine-tuned LLMs specifically to output better calibrated uncertainty estimates or to know when to say “I don’t know” [13, 28]. However, these methods require collecting labeled data or rewards and then retraining the model – which may not be feasible for very large or closed-source models. In contrast, our method introduces no new training at all: it works as an inference-time plug-in that could be applied to any pre-trained model (even black-box API models, provided one can extract a confidence measure like probabilities or logits).
@@ IV-A Uncertainty Metrics for Model Answers
These metrics are inversely related (a high $\mathrm{MSP}$ usually implies low entropy, and vice versa), but we consider both for completeness. Entropy captures the notion of hesitation in the distribution, while $\mathrm{MSP}$ directly captures confidence in the chosen answer. From a Bayesian decision perspective, if the model’s posterior over options is $P$, the Bayes-optimal 0–1 decision is $\hat{y}=\arg\max_{i}P_{i}$, and $1-\mathrm{MSP}(P)$ upper-bounds the posterior risk of choosing $\hat{y}$ under the model’s own beliefs. Motivated by calibration theory, we also define two proper scoring rules [45], which we report as secondary diagnostics in Section VI-A:
@@
To test this hypothesis, we ran DeepSeek-v3.2 on a large set of multiple-choice questions and extracted the model’s answer distribution over the provided options from token log-probabilities (converted to normalized probabilities). We then computed $H$ and Max $P$ for each question and compared these uncertainty metrics between items that were answered correctly vs. incorrectly post hoc.
@@ IV-B Empirical Trigger Criterion
IV-C Reinforcement Prompt Design
A critical component of Reinforcement Inference is the design of the second-pass prompt. The goal is to make the model aware that its prior attempt was uncertain, while avoiding any task-specific hints or ground-truth leakage.
In our implementation, the second-pass input consists of three parts:
@@ IV-C Reinforcement Prompt Design
    Original question block: the same question and options in the same structure as the first pass (e.g., Question: ..., Options: ..., Answer: ...).
@@
Importantly, this reprompting does not tell the model that it was wrong and does not provide the correct option. It only provides the previous output as context and asks for a fresh derivation under an explicit uncertainty cue.
IV-D Ensuring No Ground-Truth Leakage
A fundamental design principle for Reinforcement Inference is that the model should not accidentally get the correct answer through our intervention unless it can infer it by itself. We took several precautions to uphold this:
  • No answer revelation: We never show the model the correct answer, nor do we indicate whether its first answer was correct or incorrect.
  • Second pass includes the prior output, but no label: The second-pass input includes the model’s own first-pass raw output (“Your previous answer was: …”) to provide context. However, we never provide any signal about correctness (no “wrong”, no “correct”, and no ground-truth option).
  • Standardized instruction: The re-ask instruction text is fixed across all questions. It only indicates uncertainty (high entropy) and asks the model to discard the prior answer and recompute from first principles.
With these measures, the second-pass interaction is essentially the model solving the same question anew, but under a slight cognitive “pressure” to be careful. It’s analogous to giving a student the same problem again after telling them “you seemed unsure, try to be thorough now.” They still have to solve it themselves, but they know they should double-check their logic. Importantly, our evaluation treats the second attempt as the model’s final answer; we only count an answer as correct if the model actually produces the correct option on its own in that pass.
In summary, our method functions as a kind of self-reinforcement loop: the model’s own uncertainty triggers further reasoning by the model itself. No external hints or answers are given. This maintains the purity of the test scenario (important for fair benchmarking) and ensures that any improvement in accuracy comes from the model’s existing knowledge being better utilized, rather than from us covertly feeding it answers.
@@ V-A Model and Experiment Settings
We conducted experiments with DeepSeek-v3.2, a state-of-the-art LLM released in late 2025. DeepSeek-v3.2 is an open-source model with approximately 685 billion parameters, trained on a massive corpus of diverse multilingual and technical data [29]. It has demonstrated strong performance on a variety of benchmarks [29]. We chose DeepSeek-v3.2 for its strong reasoning ability, as we wanted a model that knows a lot but can still be challenged by difficult questions – an ideal candidate to test uncertainty-based reinforcement inference. Table I summarizes the inference parameters used across all experiments.
We used the model in a zero-shot setting with deterministic greedy decoding (temperature $t=0$) so that outputs and probabilities were stable across runs. The prompt format was kept consistent across first and second passes except for the added instruction in the second pass. Specifically, we formatted each multiple-choice question as a short prompt containing the question and the labeled options (A through J), and we instructed the model to output only the letter of the answer (e.g., “A”). This follows best practices from prior work to avoid variability in answer phrasing [31]. In the second pass, the same question format was used but with the reinforcement instruction prepended at the beginning of the prompt.
To extract the model’s answer probabilities, we used DeepSeek’s API to obtain log-probabilities for the output tokens. Since the answer in our setup is a single token (one of “A” through “J”), we could directly retrieve the probability the model assigned to each option at the answer position. These probabilities were normalized to sum to 1 over the ten options, and then we computed the entropy $H$ and Max $P$ from these values as described earlier.
@@ V-B Dataset and Questions
We drew our evaluation questions from MMLU-Pro, which is an enhanced subset of the Massive Multitask Language Understanding benchmark focusing on professional or expert-level exams [32, 33]. MMLU is a collection of 57 subject areas (history, math, science, law, medicine, etc.) with questions ranging from elementary level to professional certification level [32]. The “Pro” subset specifically includes domains like medical board exams, law school exams, advanced mathematics, and other graduate-level or professional topics [35]. These questions are known to be especially challenging for LLMs – indeed, Hendrycks et al. (2021) noted that models often perform near chance on some professional subjects and, crucially, do not know when they are wrong [32, 36].
@@ V-E Statistical Tests
In addition to uncertainty metrics derived from the predictive distribution, we evaluate model confidence using proper scoring rules, which assess the quality of probabilistic predictions with respect to the ground-truth label. Proper scoring rules are designed such that the expected score is minimized when the predicted distribution matches the true data-generating distribution, thereby encouraging both accuracy and calibration.
We consider two widely used proper scoring rules: the Brier score and the negative log-likelihood (NLL). The Brier score measures the squared error between the predicted probability vector and the one-hot encoded ground-truth label, penalizing both overconfident incorrect predictions and underconfident correct predictions. The negative log-likelihood corresponds to the cross-entropy loss evaluated at the true label and grows unbounded as the predicted probability of the correct option approaches zero, making it particularly sensitive to extreme miscalibration. Together, these metrics provide complementary views of probabilistic quality: while entropy and maximum probability characterize intrinsic model uncertainty at inference time, Brier score and NLL quantify how well the model’s confidence aligns with empirical correctness.
@@ VI Results
We first analyze the baseline one-shot pass and ask whether the model’s own uncertainty correlates with correctness. We find a clear separation: DeepSeek-v3.2 is typically more uncertain on questions it answers incorrectly. Figure 2 summarizes this baseline relationship between uncertainty (entropy and maximum probability) and correctness.
@@ VI-B Impact of Reinforcement Inference on Accuracy (Targeted Re-asking, TR)
Table IV reports the behavior of proper scoring rules (Brier score and negative log-likelihood, NLL) before and after uncertainty-triggered re-asking. While targeted re-asking (TR) substantially improves overall accuracy, we observe that for the remaining incorrect answers, both Brier score and NLL increase after TR. This effect is expected and reflects a change in the conditional distribution of errors rather than a degradation in model quality. Specifically, TR preferentially corrects errors associated with moderate uncertainty, leaving a residual set of incorrect instances that are intrinsically harder and for which the model assigns extremely low probability to the true option. As a result, the average loss under proper scoring rules—which strongly penalize near-zero probability assigned to the ground-truth label—increases for the incorrect subset after re-asking.
Overall, these results demonstrate that Reinforcement Inference succeeded in translating uncertainty awareness into concrete accuracy gains at scale. The model effectively self-corrected: when prompted about its own uncertainty, it was able to correctly solve a large fraction of the problems it initially missed. This suggests that the knowledge needed to answer many of those questions was latent in the model, and the second-pass prompt helped unlock it.
@@ VI-C Ablation Study: Uniform Re-asking (UR)
Critically, this condition does not include (i) the model’s first-round prediction, nor (ii) any actual uncertainty signal derived from a prior output, since no initial pass is performed. Thus, this ablation controls for both answer feedback and selective triggering, leaving only the effect of the re-asking instruction and CoT encouragement.
@@
Implications. This result rules out the hypothesis that our improvements stem from the re-asking prompt or generic CoT elicitation alone. Instead, it provides evidence that the performance gains are driven by the interaction between (a) conditioning on the model’s own first-round answer and (b) selectively invoking re-asking based on a measured uncertainty signal. In the absence of a concrete prior attempt and a grounded justification for re-asking, the same prompt fails to yield benefits. This supports our central claim that uncertainty-aware, answer-conditioned re-asking, rather than prompt-level encouragement to “think harder,” is the key factor underlying the observed improvements.

### v3

@@ I Introduction
Despite the remarkable performance, LLMs can produce confidently incorrect answers. A well-known challenge is that models often lack calibrated confidence – they may not recognize when their answers are likely wrong [1, 2]. This can lead to misleading outputs in high-stakes settings (for example, an LLM hallucinating a plausible-sounding but false explanation with full certainty [3]). If an AI could know when it is unsure, it might mitigate errors by double-checking or deferring answers. Indeed, prior studies on tasks like multiple-choice QA have noted that an LLM’s softmax probabilities carry information about correctness [2, 4]. Intuitively, a model that assigns one option nearly 100% probability is likely confident (and often correct), whereas if it spreads probability roughly uniformly across all options, it is essentially guessing [2]. Recent evaluations confirm that wrong answers tend to have lower maximum softmax probability (MSP) than correct answers in well-trained models [2, 5]. In short, uncertainty correlates with mistakes.
How can we use this insight to improve model reasoning? One approach is to have the model abstain or refuse to answer when uncertainty is high [6]. Refusal might be safer in open-ended applications, but in many scenarios we do want the model to attempt an answer – just with more care if it was initially unsure. Humans do this naturally: if unsure of an answer, a person might pause and double-check their reasoning or try a different approach before responding [7]. This inspires the question: Can an LLM similarly “think twice” when it hesitates, to produce a more accurate answer, all at inference time?
@@
In this paper, we introduce Reinforcement Inference, an inference-time framework that lets an LLM use its own uncertainty to decide when to “think twice.” We run a deterministic first pass, monitor entropy and maximum softmax probability (MSP), and trigger a single additional, more deliberate attempt only when the model is uncertain. Crucially, this uses no labels or external feedback and requires no retraining.
We argue that default one-shot greedy decoding can understate a fixed model’s true capability: in deterministic deployments, many failures reflect premature commitment under ambiguity rather than missing knowledge [7]. Reinforcement Inference operationalizes this view by treating uncertainty as a first-class control signal, yielding an adaptive compute policy with predictable behavior on easy cases and extra deliberation only when needed.
More broadly, our results motivate an entropy-aware training direction that explicitly shapes correctness–confidence alignment. The resulting “capability gap” between one-pass inference and uncertainty-conditioned deliberation can serve as a practical diagnostic for system design and model optimization.
@@ III-A Confidence and Uncertainty in LLMs
Recent work studies how well an LLM’s predictive probabilities reflect correctness and whether models can explicitly reason about their own confidence. Maximum softmax probability (MSP) is a common proxy and tends to be lower for incorrect and out-of-distribution inputs [1]. Studies show that softmax distributions can predict correctness in multiple-choice QA and can be used for selective abstention [2, 6]. Related work studies Bayesian/ensemble uncertainty estimation and calibration via fine-tuning [5, 8]. We likewise use the model’s output distribution as an uncertainty signal, but rather than training calibration or abstention, we trigger self-correction at inference time.
@@ III-B Chain-of-Thought and Self-Consistency
Chain-of-thought (CoT) prompting can improve performance by eliciting intermediate reasoning [9]. Self-consistency extends CoT by sampling multiple reasoning paths and selecting the majority answer [10], yielding sizable gains on tasks such as math and commonsense reasoning. However, it is compute-intensive and uses multiple attempts regardless of confidence. In contrast, Reinforcement Inference triggers only a single additional pass and does so selectively when the first pass is uncertain, saving compute on easy queries.
@@ III-C Verifier Models and Answer Re-Ranking
Another line improves correctness via explicit verification. Some methods train separate verifiers to accept/reject model answers, while others prompt the model to verify itself: Chain-of-Verification (CoVe) iteratively generates an answer, asks and answers follow-up checks, and revises the output, reducing hallucinations [11]. Related self-critique approaches similarly analyze and refine the initial output. Compared to these multi-step pipelines, Reinforcement Inference is lightweight: a single uncertainty-triggered second pass without learned verifiers or explicit verification questions.
@@ III-D Supervised Retraining for Reasoning
A common way to improve accuracy or calibration is additional training or fine-tuning on reasoning data or aligning with RLHF. Recent work has also fine-tuned LLMs to output better-calibrated uncertainty or to abstain (“I don’t know”) [12]. However, such approaches require labeled data/rewards and retraining, which can be impractical for very large or closed-source models. In contrast, our method requires no training: it is an inference-time plug-in applicable to any model where a confidence signal (e.g., probabilities/logits) can be obtained.
@@ IV-A Uncertainty Metrics for Model Answers
These metrics are inversely related (a high $\mathrm{MSP}$ usually implies low entropy, and vice versa), but we consider both for completeness. Entropy captures the notion of hesitation in the distribution, while $\mathrm{MSP}$ directly captures confidence in the chosen answer. From a Bayesian decision perspective, if the model’s posterior over options is $P$, the Bayes-optimal 0–1 decision is $\hat{y}=\arg\max_{i}P_{i}$, and $1-\mathrm{MSP}(P)$ upper-bounds the posterior risk of choosing $\hat{y}$ under the model’s own beliefs. Inspired by classical post-hoc calibration methods such as Platt scaling [13], we also define two proper scoring rules [14], which we report as secondary diagnostics in Section VI-A:
@@
To test this hypothesis, we ran DeepSeek-v3.2 on a large set of multiple-choice questions and extracted the model’s answer distribution over the provided options from token log-probabilities (converted to normalized probabilities). We then computed $H$ and $\mathrm{MSP}$ for each question and compared these uncertainty metrics between items that were answered correctly vs. incorrectly post hoc.
@@ IV-B Empirical Trigger Criterion
IV-C Second-Pass Prompting Protocol
A critical component of Reinforcement Inference is the design of the second-pass prompt. The goal is to make the model aware that its prior attempt was uncertain, while avoiding any task-specific hints. In our implementation, the second-pass input consists of three parts:
@@ IV-C Second-Pass Prompting Protocol
    Original question block: the same question and options in the same structure as the first pass (e.g., Question: ..., Options: ...).
@@
Importantly, the reprompt never tells the model it was wrong and never reveals the correct option. We only include the first-pass output as context and request a fresh derivation under an explicit uncertainty cue, so the second pass re-solves the same question with added pressure to be careful (analogous to asking a student to redo a problem after noting hesitation).
In summary, this is a self-reinforcement loop: uncertainty triggers additional reasoning without external hints, preserving a clean evaluation and attributing gains to better use of the model’s existing knowledge rather than information leakage.
@@ V-A Model and Experiment Settings
We conducted experiments with DeepSeek-v3.2, a state-of-the-art LLM released in late 2025. It is an open-source model with approximately 685 billion parameters, trained on a massive multilingual and technical corpus and strong across benchmarks [15]. We chose it as a strong yet challenging testbed for uncertainty-based Reinforcement Inference. Table I summarizes the inference parameters used across all experiments.
We evaluate in a zero-shot, deterministic setting with greedy decoding (temperature $t=0$) to ensure stable outputs and probabilities. Both passes use the same multiple-choice format (question + options A–J) and prompt the model to output answering process and option letter [16]; the second pass simply prepends the reinforcement instruction.
To obtain answer probabilities, we query the API for token-level log-probabilities. Because the answer is a single token (A–J), we directly read out the probability of each option, normalize over the ten choices, and compute the entropy $H$ and $MSP$ as described earlier.
@@ V-B Dataset and Questions
We drew our evaluation questions from MMLU-Pro, which is an enhanced subset of the Massive Multitask Language Understanding benchmark focusing on professional or expert-level exams [17, 18]. MMLU is a collection of 57 subject areas (history, math, science, law, medicine, etc.) with questions ranging from elementary level to professional certification level. The “Pro” subset specifically includes domains like medical board exams, law school exams, advanced mathematics, and other graduate-level or professional topics. These questions are known to be especially challenging for LLMs – models often perform near chance on some professional subjects and, crucially, do not know when they are wrong [19].
@@ V-E Statistical Tests
In addition to uncertainty metrics from the predictive distribution, we assess confidence using proper scoring rules, which evaluate probabilistic predictions against the ground-truth label and are minimized in expectation by the true distribution.
We use the Brier score and negative log-likelihood (NLL). The Brier score is the squared error between the predicted probability vector and the one-hot label, while NLL (cross-entropy at the true label) strongly penalizes assigning low probability to the correct option. These complement entropy/MSP by measuring how well confidence aligns with empirical correctness.
@@ VI Results
We first analyze the baseline one-shot pass and ask whether the model’s own uncertainty correlates with correctness. We find a clear separation: DeepSeek-v3.2 is typically more uncertain on questions it answers incorrectly. Figure 2 summarizes this baseline relationship between uncertainty (entropy and maximum probability) and correctness [15].
@@ VI-B Impact of Reinforcement Inference on Accuracy (Targeted Re-asking, TR)
Table IV reports proper scoring rules (Brier score and negative log-likelihood, NLL) before and after uncertainty-triggered re-asking. Although targeted re-asking (TR) improves accuracy, Brier and NLL can increase on the remaining incorrect cases. This is expected: TR mainly fixes moderate-uncertainty mistakes, leaving a smaller set of intrinsically hard errors where the model assigns near-zero probability to the true option, which these losses penalize strongly.
Overall, Reinforcement Inference translates uncertainty awareness into large accuracy gains via effective self-correction, suggesting that many initially missed answers are latent and can be recovered with a second-pass prompt.
@@ VI-C Ablation Study: Uniform Re-asking (UR)
This condition does not include (i) the model’s first-round prediction, nor (ii) any actual uncertainty signal derived from a prior output, since no initial pass is performed. Thus, this ablation controls for both answer feedback and selective triggering, leaving only the effect of the re-asking instruction and CoT encouragement.
@@
Implications. This indicates that gains do not come from the prompt or generic CoT elicitation alone; they require uncertainty-triggered, selective re-asking conditioned on a concrete first-round attempt. In short, the key is uncertainty-aware, answer-conditioned re-asking rather than urging the model to “think harder.”